# Olaph: Improving Factuality in Biomedical Long-form Question Answering

### v1

@@ 4 How to Train Olaph?
We introduce Olaph (Optimizing Large language models’ Answers with Preferences of mitigating Hallucination), a simple and novel framework designed to optimize responses of language models (LMs) by aligning them with preference collections. We first train with supervised fine-tuning (SFT) to familiarize the model with the question-answering task using relatively small data samples (Section 4.1). Then, we obtain $k$ predictions sampled with temperature scaling [32] (Section 4.2). We evaluate these predictions using diverse evaluation metrics to gauge the LLM’s generation ability. Then, we make a preference set using the best sample by computing the score of evaluation metrics and train with direct preference optimization (DPO) [33] (Section 4.3). Finally, we iteratively train LLMs through DPO until convergence (Section 4.4).
@@ 4.1 Supervised Fine-tuning
SFT leverages relatively smaller data of labeled samples to tailor a pre-trained LLM to specific tasks [31, 58]. Rather than training on human annotations or pseudo-optimal responses generated by larger language models, we set a self-generated response as a labeled answer to remove the dependency on resources in annotation datasets [59, 60]. We train the LLM with SFT as below,
@@
4.2 Automatic Evaluation of Machine-Generated Texts
We depict the overall procedure in Figure 3. After initializing with $\pi_{SFT}$, we obtain sampled predictions through temperature scaling [32]. We generate $k$ predictions (we use $k$=6 here): one for deterministic prediction and five for sampling predictions. We then sort all sampled predictions with the following weighted sum score of the automatic evaluation criteria,
@@ 4.2 Automatic Evaluation of Machine-Generated Texts
where $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ reflect the weighted importance of each evaluation metric set as hyperparameters respectively. $r_{1},r_{2},r_{l}$ refer to Rouge-score [38] that measures how much similar words are used. BL and BS refer to BLEURT [41] and BERTScore [42] which are used to measure semantic similarity. HL and CP refer to Hallucination and Comprehensiveness which are used to measure inclusion of crucial claims [22]. We deduct the HL score in the evaluation metric because this is the only score that affects to language model’s response to get worse.
We sort $k$ sampled predictions with based on the score of the weighted sum of evaluation metrics in Equation 4. Then, we use a pre-determined threshold to distinguish preferences and create the preference set (high score) and the dispreference set (low score) to guide how language models should respond. We describe details of training through the preference set in the following section.
Figure 3: Overall Olaph framework. We iteratively implement the following steps to train LLMs. If a patient asks a question about the details of Lexapro, we generate $k$ predictions with temperature scaling. These predictions are evaluated based on three main categories of evaluation metrics. We compute the weighted sum of the categories and sort predictions with the score. We distinguish two sets (preferred and dispreferred) using a pre-determined threshold to train the language model through preference optimization [33]. Finally, we obtain the preferred answer to the patient’s question. Here, we omit the SFT training part.
@@ 4.3 Direct Preference Optimization
We use the concept of direct preference optimization (DPO) [33] to optimize a student model $\pi_{\theta}$ to maximize the likelihood of generating less hallucinated text [46, 47, 48, 50]. We agree with the notion that language models already embody a certain level of knowledge about potential responses [61, 62, 63]. Hence, we believe that among the responses generated through sampling, there may be predictions that closely resemble the desired ways of answering the question. Therefore, we aim to enhance the quality of long-text responses through DPO learning, adjusting the student model $\pi_{\theta}$ finely to generate the preferred response $r_{w}$ over the dispreferred response $r_{l}$. We train the student model $\pi_{\theta}$ as below,
@@
$r_{\theta}(x,r)$ $=\beta(\text{log}~{}\pi_{\theta}(r|x)-~{}\text{log}~{}\pi_{SFT}~{}(r|x))$
@@
4.4 Iterative Learning of LLMs via DPO
Our Olaph framework iteratively trains LLMs through DPO multiple times, presenting a situation where distinguishing between preferred and dispreferred responses becomes increasingly challenging. Through this process, we have two benefits: (1) Generating a synthetic preference set step-by-step eliminates dependency on human-annotated datasets, which require labor-intensive work. (2) Applying multifaceted evaluation metrics enhances the overall quality of long-form answers, showing improvement in unseen evaluation metrics as well. These benefits lead us to design Olaph framework to train iteratively until convergence.
In summary, our Olaph framework utilizes automatic evaluation to select the best response in sampling predictions and answers questions in our preferred manner, training the model step-by-step to reduce hallucinations and include crucial medical claims.
@@ 5 Experimental Settings
Data consisting of questions that patients or customers frequently inquire about in the biomedical or clinical domain is very scarce. Furthermore, there is a lack of collected training datasets for LFQA benchmarks, making it difficult to assess the effectiveness of our Olaph framework. Therefore, we leave one dataset as a test set and train on the remaining datasets. For example, in Table 2, we evaluate LiveQA [34] dataset while training on the MedicationQA [35], HealthSearchQA [1], and K-QA [22] datasets. We provide further details of training and inference settings in Appendix A.2.
@@ 6 Experiments & Analysis
In this section, we first explore the generation ability of the large language models (LLMs) using the reconstructed MedLFQA. Then, we describe the observations after applying our Olaph framework to mitigate hallucinations.
We have three research questions as follows: (1) How well can open-foundation LLMs answer clinical questions? (2) How many steps of iterative learning are necessary to enhance the generation ability of 7B language models, up to that of GPT-4? (3) Do the results align with other factuality metrics such as FActScore [20], which are not used in our fine-grained evaluation metrics?
@@
We perform a zero-shot evaluation to assume the real scenarios where users utilize LLMs. We provide the overall results in Table 2. We observe that base foundation models such as LLaMA2 [64] and Mistral [65] answer properly on some datasets but not consistently. The responses of these models show lower factuality (low Comprehensiveness and Hallucination) while preserving the score of words composition and semantic similarity.
Three biomedical language models that underwent instruction tuning exhibit different patterns compared to the base models. Two of the models, Meditron [2] and BioMistral [6], which were trained on instructions related to the biomedical or clinical domain, record very low scores in terms of factuality. The results indicate that given a medical question, the answers are composed of hallucinated responses with less crucial claims. However, Self-BioRAG [7], which was trained on diverse instructions containing long-form question answering, consistently performs well in providing answers to medical questions.
@@
This analysis aims to investigate the extent to which the ability of long-text generation can be enhanced through iterative learning. We conduct the analysis using the K-QA golden dataset [22] which contains answers annotated by medical experts. We depict the performance improvements in Figure 4. We represent the median of the three evaluation metrics used to select the preferred response as lines. The underlying colors represent the lower and upper bounds of the model for each step. Since the answers were annotated using GPT-4 API calls, we set the upper bound for long-form answers generated by GPT-4 when solving the K-QA golden dataset.
In the initial step (Step 0), the BioMistral model shows low scores for all evaluation metrics selected. As the steps progressed, the performance improved and approached the scores of GPT-4 response. We find that performance tends to saturate after DPO (Step 2) training. Finally, after iterative DPO training (Step 3), we observe that the 7B model reaches the upper bound performance. Experiment results for other 7B models can be found in Appendix A.3.
@@
Our fundamental inquiry revolves around whether the responses generated by the LLM trained with our Olaph framework have indeed improved in terms of factuality. To ascertain this, our focus is on evaluating the degree to which factuality has increased based on the FActScore [20] metric, which is not used during training.
We depict the FActScore performance at each step in Figure 5. FActScore involves the selection of context-containing pages from topics chosen from Wikipedia dump. Subsequently, the generated responses are segmented into atomic facts, and GPT-3.5 is employed to confirm whether the identified context supports the atomic facts. In the case of the K-QA golden dataset [22], as no topic is provided, we select topics from a set of entities within the questions and measure the FActScore to ensure connections to appropriate pages. Additionally, considering the potential lack of biomedical knowledge in the Wikipedia dump, we further construct the domain-specific knowledge following Self-BioRAG [7]. The biomedical knowledge consists of four source data: PubMed Abstract, PMC Full-text, Clinical Guidelines, and English Medical Textbooks. We use a domain-specific retriever MedCPT [66] off-the-shelf to retrieve the relevant document. We provide the details of the knowledge source and retriever in Appendix A.4.
@@
Refer to captionRefer to caption
Figure 5: We evaluate factuality using FActScore performance which is not used evaluation metric during training. We report FActScore without length penalty as a metric. We supply domain-specific knowledge due to the potential lack of biomedical knowledge. We also provide the GPT-4 score for the upper bound of FActScore performance.
@@ 7 Conclusion, Limitations, and Future Work
We introduce Olaph, a novel framework that reduces hallucinations and includes crucial claims by utilizing automatic evaluation to select the best response in sampling predictions and designing to answer questions in our preferred manner. We also introduce MedLFQA which is reconstructed into a unified format containing long-form answers and crucial statements to facilitate automatic evaluation. We demonstrate that current 7B LLMs are not reliable enough to answer medical questions in the long-text format. Learning through our Olaph framework step-by-step, consisting of SFT and iterative DPO learning, 7B models can generate long answers that contain enough factuality, semantic similarities, and word composition.
One limitation could be that false knowledge or hallucinations may exist in the reconstructed LFQA benchmark answers and statements. Since it is beyond our capabilities to correct them manually. This issue could be addressed by involving medical experts in the review process or by finding better ways to generate them. Furthermore, we compare and analyze models with a size of 7B parameters, which is suitable for environments with limited resources. It is necessary to consider models with smaller or larger parameter sizes to determine the effectiveness of our method in confirming results and analysis. However, if the model is smaller than 7B, there is a lower probability of generating correct predictions, and sampling predictions may not yield proper responses. Also, MedLFQA consists of biomedical knowledge predefined within a fixed timestamp, which could raise outdated issues in the future. Finally, there is a possibility of error propagation in the evaluation due to the use of trained NLI models. However, our approach is aimed at establishing evaluation metrics that can replace tremendous API costs in a cost-effective manner.
With the advancement of LLM’s generation abilities, our study demonstrates that 7B LLMs are capable of producing long-text medical answers at a desirable level, when trained with the appropriate data and methods. For future work, if 7B or even larger LLMs are enabled to comprehend a patient’s history and engage in multi-turn conversations, they could be sufficiently utilized to assist physicians as a conversation agent.
Acknowledgements
@@ A.1 Criteria of Generated Answer Evaluation
We follow the criteria for evaluating GPT-4 generated answers used in the previous works [1]. The authors provide nine criteria to evaluate language models’ responses in a fine-grained manner. We provide the details in the Table 3. The four criteria given above are preferable when selected (row 1-4), while the five criteria given below indicate a better answer when not selected (row 5-9). We use these criteria on pairwise evaluation between GPT-4 and medical expert answers.
Table 3: Nine Criteria used for pairwise evaluation.
Criteria Definition
Alignment with Medical Consensus (MC) Which answer better reflects the current consensus of the scientific and clinical community?
@@
Omission of Important Information (OII) Which answer omits more important information?
@@
A.2 Experimental Details

### v2

@@ 4 How to Train Olaph?
We introduce Olaph (Optimizing Large language models’ Answers with Preferences of mitigating Hallucination), a simple and novel framework designed to optimize responses of language models (LMs) by aligning them with preference collections. We first train with supervised fine-tuning (SFT) to familiarize the model with the question-answering task using relatively small data samples (Section 4.1). Then, we obtain $k$ sampled predictions sampled with temperature sampling (Guo et al., 2017). We evaluate these predictions using diverse evaluation metrics to distinguish preferred and dispreferred answer (Section 4.2). Then, we make a preference set in every steps using previous-step model with self-generated responses and train with direct preference optimization (DPO) (Rafailov et al., 2024) (Section 4.3). Finally, we iteratively tune our LLMs until convergence (Section 4.4).
@@ 4.1 Supervised Fine-tuning
SFT leverages relatively smaller data of labeled samples to tailor a pre-trained LLM to specific tasks (Ouyang et al., 2022; Yu et al., 2023). Rather than training on human annotations or pseudo-optimal responses generated by larger language models, we set a self-generated response as a labeled answer of next step training to remove the dependency on resources in annotation datasets (Chen et al., 2024; Wu et al., 2024). In other words, we generate multiple self-generated responses using sampling-based inferences of temperature sampling, and from these responses we select the one that scores the highest according to the automatic evaluation categories as the gold-standard label for the next step of training. We train the LLM with SFT as below,
@@
4.2 Cost-effective and Multifacted Automatic Evaluation
We depict the overall procedure in Figure 3. After initializing with $\pi_{SFT}$, we obtain sampled predictions through temperature sampling (Guo et al., 2017). We generate $k$ predictions (we use $k$=6 here): one for deterministic prediction and five for sampling predictions. We then sort all sampled predictions with the following weighted sum score of the automatic evaluation criteria,
@@ 4.2 Cost-effective and Multifacted Automatic Evaluation
where $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ reflect the weighted importance of each evaluation metric set as hyperparameters respectively. $r_{1},r_{2},r_{l}$ refer to Rouge-score (Lin, 2004) that measures how much similar words are used. BL and BS refer to BLEURT (Sellam et al., 2020) and BERTScore (Zhang et al., 2019) which are used to measure semantic similarity. HL and CP refer to Hallucination and Comprehensiveness which are used to measure inclusion of crucial claims (Manes et al., 2024). We deduct the HL score in the evaluation metric because this is the only score that affects to language model’s response to get worse.
We sort $k$ sampled predictions with based on the score of the weighted sum of evaluation metrics in Equation 4. Then, we use a pre-determined threshold to distinguish preferences and create the preference set (high score) and the dispreference set (low score) to guide how language models should respond. [^6] We provide the sensitivity analysis of our hyperparameters ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$, and pre-determined threshold) in Appendix A.2. We describe details of training through the preference set in the following section.
Figure 3: Overall Olaph framework. We iteratively implement the following steps to train LLMs (Step 1-4). If a patient asks a question about the details of Lexapro, we generate $k$ predictions with temperature sampling (Step 1). These predictions are evaluated based on three main categories of our preferred evaluation metrics. We compute the multifaceted automatic evaluation and sort predictions with the score (Step 2). We distinguish two sets (preferred and dispreferred) using a pre-determined threshold to construct the synthetic alignment pair dataset (Step 3). We then train the LLMs through preference optimization such as DPO (Rafailov et al., 2024) (Step 4). Finally, we obtain the preferred answer to the patient’s question. Here, we omit the SFT training part.
@@ 4.3 Direct Preference Optimization
We use the concept of direct preference optimization (DPO) (Rafailov et al., 2024) to optimize a student model $\pi_{\theta}$ to maximize the likelihood of generating less hallucinated text (Tian et al., 2023; Zhang et al., 2023; Kang et al., 2024; Dahl et al., 2024). We agree with the notion that language models already embody a certain level of knowledge about potential responses (Saunders et al., 2022; Kadavath et al., 2022; Li et al., 2024b). Hence, we believe that among the responses generated through sampling, there may be predictions that closely resemble the desired ways of answering the question. Therefore, we aim to enhance the quality of long-text responses through DPO learning, adjusting the student model $\pi_{\theta}$ finely to generate the preferred response $r_{w}$ over the dispreferred response $r_{l}$. We train the student model $\pi_{\theta}$ as below,
@@
$\vspace{0.1cm}r_{\theta}(x,r)$ $=\beta(\text{log}~{}\pi_{\theta}(r|x)-~{}\text{log}~{}\pi_{SFT}~{}(r|x))$
@@
4.4 Iterative Learning with Self-generated Preference Set
Our Olaph framework iteratively trains LLMs through DPO multiple times, presenting situations where each step contains distinguishing between preferred and dispreferred responses based on the cost-effective automatic evaluations to make preference set. Through this process, we have two benefits: (1) In each step, constructing a synthetic preference set with self-generated responses using temperature sampling can eliminate dependency on human-annotated datasets, which require labor-intensive work. (2) Applying cost-effective and multifaceted evaluation metrics enhances the overall quality of long-form answers, showing improvement in unseen evaluation metrics as well. These benefits lead us to design Olaph framework to train iteratively until convergence.
In summary, our Olaph framework utilizes cost-effective and multifaceted automatic evaluation to construct a synthetic preference set and answers questions in our preferred manner, which leads us to train LLMs step-by-step to reduce hallucinations and include crucial medical claims.
@@ 5 Experimental Settings
Data consisting of questions that patients or customers frequently inquire about in the biomedical or clinical domain is very scarce. All five datasets comprising MedLFQA consist only of test datasets, with no separate train datasets. Therefore, there is a lack of collected training datasets to evaluate these benchmarks, making it difficult to assess the effectiveness of our Olaph framework. In this situation, we designated one dataset as the test dataset and used the remaining datasets as the train datasets for training purposes. In other words, we leave one dataset as a test set and train on the remaining datasets same concept as a cross-validation. For example, we evaluate LiveQA dataset while training on the MedicationQA, HealthSearchQA, and K-QA datasets (row 1 in Table 2). If we want to evaluate HealthSearchQA dataset, then we train with LiveQA, MedicationQA, and K-QA datasets (row 3 in Table 2). We provide further details of training and inference settings in Appendix A.3.
@@ 6 Experiments & Analysis
In this section, we first explore the generation ability of the large language models (LLMs) using the reconstructed MedLFQA dataset. Then, we describe the observations after applying our Olaph framework to mitigate hallucinations. Thus, we have three research questions as follows: (1) How well can open-foundation and proprietary LLMs answer clinical questions? (2) How many steps of iterative learning are necessary to enhance the generation ability of 7B language models, up to that of GPT-4? (3) Do the results align with other factuality metrics such as FActScore (Min et al., 2023), which are not used in our fine-grained evaluation metrics?
@@
We perform a zero-shot evaluation to assume the real scenarios where users utilize LLMs. We provide the overall results in Table 2. We observe that base foundation models such as LLaMA2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023) answer properly on some datasets but not consistently. The responses of these models show lower factuality (low Comprehensiveness and Hallucination) while preserving the score of words composition and semantic similarity.
Three biomedical language models that underwent instruction tuning exhibit different patterns compared to the base models. Two of the models, Meditron (Chen et al., 2023) and BioMistral (Labrak et al., 2024), which were trained on instructions related to the biomedical or clinical domain, record very low scores in terms of factuality. The results indicate that given a medical question, the answers are composed of hallucinated responses with less crucial claims. However, Self-BioRAG (Jeong et al., 2024), which was trained on diverse instructions containing long-form question answering, consistently performs well in providing answers to medical questions.
Additionally, we use three proprietary LLMs to answer the clinical questions in Table 3. In our observation, proprietary LLMs perform remarkably well in generating long-form responses to clinical questions compared to the open-foundation models. However, researchers cannot reach to these black-box LLMs to elicit and update their knowledge through training. Thus, we try to focus our Olaph approach on low resource (under 7B) and open-source models in the following sections.
Table 3: We use MedLFQA to evaluate three proprietary language models. The evaluation metrics are composed of three in total: words composition, semantic similarity, and factuality. The numbers are the results obtained by zero-shot experiments asking the question as prompt into the LLMs.
MedLFQA Dataset Evaluation Metrics Proprietary LLMs
GPT-3.5-Turbo Claude 3 Sonnet GPT-4o
LiveQA Words Composition 36.6 44.3 48.5
Semantic Similarity 108.0 116.5 75.3
Factuality 55.4 71.2 75.3
MedicationQA Words Composition 38.2 48.9 50.0
Semantic Similarity 109.8 122.3 121.2
Factuality 58.3 79.9 81.2
HealthSearchQA Words Composition 29.7 41.2 39.7
Semantic Similarity 105.3 115.1 112.3
Factuality 48.0 71.3 65.6
K-QA Golden Words Composition 35.6 48.5 51.7
Semantic Similarity 109.7 119.3 122.1
Factuality 52.5 82.8 85.9
K-QA Silver Words Composition 36.2 51.3 52.9
Semantic Similarity 112.0 117.7 119.5
Factuality 51.3 80.1 83.7
@@
This analysis aims to investigate the extent to which the ability of long-text generation can be enhanced through iterative learning. We conduct the analysis using the K-QA golden dataset (Manes et al., 2024) which contains answers annotated by medical experts. We depict the performance improvements in Figure 4. We represent the median of the three evaluation metrics used to select the preferred response as lines. The underlying colors represent the lower and upper bounds of the model for each step. Since the answers were annotated using GPT-4 API calls, we set the upper bound for long-form answers generated by GPT-4 when solving the K-QA golden dataset.
In the initial step (Step 0), the BioMistral model shows low scores for all evaluation metrics selected. As the steps progressed, the performance improved and approached the scores of GPT-4 response. We find that performance tends to saturate after DPO (Step 2) training. Finally, after iterative DPO training (Step 3), we observe that the 7B model reaches the upper bound performance. We provide other results of 7B models in Appendix A.4.
@@
Our fundamental inquiry revolves around whether the responses generated by the LLM trained with our Olaph framework have indeed improved in terms of factuality. To ascertain this, our focus is on evaluating the degree to which factuality has increased based on the FActScore (Min et al., 2023) metric, which is not used during training.
We depict the FActScore performance at each step in Figure 5. FActScore involves the selection of context-containing pages from topics chosen from Wikipedia dump. Subsequently, the generated responses are segmented into atomic facts, and GPT-3.5 is employed to confirm whether the identified context supports the atomic facts. In the case of the K-QA golden dataset (Manes et al., 2024), as no topic is provided, we select topics from a set of entities within the questions and measure the FActScore to ensure connections to appropriate pages. Additionally, considering the potential lack of biomedical knowledge in the Wikipedia dump, we further construct the domain-specific knowledge following Self-BioRAG (Jeong et al., 2024). The biomedical knowledge consists of four source data: PubMed Abstract, PMC Full-text, Clinical Guidelines, and English Medical Textbooks. We use a domain-specific retriever MedCPT (Jin et al., 2023) off-the-shelf to retrieve the relevant document. We provide the details of the knowledge source and retriever in Appendix A.5.
@@
Figure 5: We evaluate factuality using FActScore performance which is not used evaluation metric during training. We report FActScore without length penalty as a metric. We supply domain-specific knowledge due to the potential lack of biomedical knowledge. We also provide the GPT-4 score for the upper bound of FActScore performance. We observe that starting with SFT shows performance degradation, but demonstrates its highest effectiveness with iterative alignment tuning.
@@ 7 Conclusion, Limitations, and Future Work
We introduce Olaph, a novel framework designed to reduce hallucinations and include crucial claims by utilizing cost-effective and multifaceted automatic evaluation to select the best response from sampled predictions and structuring answers in a preferred format. We also present MedLFQA which has been reconstructed into a unified format containing long-form answers and crucial statements, facilitating cost-effective automatic evaluation. Our findings show that current 7B LLMs are not reliable enough to generate long-form answers to medical questions. However, by utilizing our Olaph framework, which includes step-by-step processes like SFT, preference set construction based on multifaceted automatic evaluation, and iterative alignment tuning, 7B models can produce answers with sufficient factual accuracy, semantic similarity, and coherent word composition.
One limitation could be that we compare and analyze models with a size of 7B parameters, which is suitable for environments with limited resources. It is necessary to consider models with smaller or larger parameter sizes to determine the effectiveness of our method in confirming results and analysis. However, if the model is smaller than 7B, there is a lower probability of generating correct predictions, and sampling predictions may not yield proper responses. Also, MedLFQA consists of biomedical knowledge predefined within a fixed timestamp, which could raise outdated issues in the future. Finally, there is a possibility of error propagation in the evaluation due to the use of trained NLI models. However, our approach is aimed at establishing evaluation metrics that can replace tremendous API costs in a cost-effective manner.
With the advancement of LLM’s generation abilities, our study demonstrates that 7B LLMs are capable of producing long-text medical answers at a desirable level, when trained with the appropriate data and methods. For future work, if 7B or even larger LLMs are enabled to comprehend a patient’s history and engage in multi-turn conversations, they could be sufficiently utilized to assist physicians as a conversation agent specialized at responding in a personalized situation.
Acknowledgments
@@ A.1 Criteria of Generated Answer Evaluation
We follow the criteria for evaluating GPT-4 generated answers used in the previous works (Singhal et al., 2023). The authors provide nine criteria to evaluate language models’ responses in a fine-grained manner. We provide the details in the Table 4. The four criteria given above are preferable when selected (row 1-4), while the five criteria given below indicate a better answer when not selected (row 5-9). We use these criteria on pairwise evaluation between GPT-4 and medical expert answers. We further compute the agreement by determining if at least two out of the three medical experts chose the same answer. Note that they have shown the high level of agreement that GPT response is highly usable.
Table 4: Nine Criteria used for pairwise evaluation. We observe an extreme level of agreement that GPT-4 response is highly available among the medical experts across all items.
Criteria Definition Agreement
Alignment with Medical Consensus (MC) Which answer better reflects the current consensus of the scientific and clinical community? 99%
@@
99%
Omission of Important Information (OII) Which answer omits more important information? 99%
@@
99%
A.2 Sensitivity Analysis of Hyperparameters
Sensitivity Analysis of $\alpha_{3}$.
We aimed to conduct a comprehensive sensitivity analysis on the hyperparameter settings to determine the factuality. In Table 5 and 6, we provide the detail experiments. In our experiments, we fixed $\alpha_{1}$ and $\alpha_{2}$ at a value of 1, while varying $\alpha_{3}$ in increments of 0.2 from 0 to 1. These experiments were carried out using the BioMistral-7B and Self-BioRAG 7B models, with training data of LiveQA, MedicationQA, HealthSearchQA, and KQA-Silver, and evaluation data of KQA-Golden dataset.
A notable observation is that while performance on evaluation metrics such as word composition and semantic similarity consistently improves, setting $\alpha_{3}$ to 0 results in minimal changes in factuality. Furthermore, increasing the value of $\alpha_{3}$ (moving to higher values) correlates with improved factuality scores. We also found that iterative DPO training reduces the standard deviation in overall scores, indicating that as training progresses, the model’s confidence increases, leading to more reliable answers.
Sensitivity Analysis of Pre-determined Threshold.
The criteria for dividing the actual preference set and dispreference set are determined according to Equation 4. The threshold defines what response should align to preferred or dispreferred response. If the model’s response exceeds the threshold, that response is included in the preferred set, while responses below the threshold are included in the dispreferred set, creating pairs that are used in next-step DPO training. In other words, the threshold helps construct the pair dataset by distinguishing between preferred and dispreferred responses based on whether their automatic evaluation scores are above or below the threshold.
To comprehensively understand why the threshold is needed, we first introduce our range of each evaluation category used in the Equation 4. We want to note that the scaling normalization was applied to ensure all evaluation metrics have an equal scale range. For example, in the case of ROUGE scores, each score value is set between 0 and 1. However, for comprehensiveness or hallucination, which measures factuality, the score range is from -100 to 100. To eliminate variations due to these scaling differences, we performed scale normalization so that ROUGE, BLEURT, and BERTScore all operate on the same scale. Below are the lower and upper bounds for each evaluation metric, followed by the score ranges for each evaluation category:
  • Words Composition: [0, 3] with $\alpha_{1}=1.0$; scaling up to [0, 300] to match the number range to the other hyperparameters.
  • Semantic Similarity: [Unknown, 2] with $\alpha_{2}=1.0$; scaling up to [Unknown, 200] to match the number range to the other hyperparameters.
  • Factuality: [-100, 100] with $\alpha_{3}=1.0$.
The pre-determined threshold of Equation 4 is set as 200. This value was intuitively determined by manually evaluating the model’s responses according to the authors’ preferences. Generally, when evaluating multiple answers, responses that scored high on average across all items exceeded 200, so this value was set as our threshold. However, recognizing that this method requires careful review, we conduct the experiments by setting a broad range of thresholds (0, 50, 100, 150, 200) in Table 7.
Table 5: BioMistral 7B performance of sensitivity analysis ($\alpha_{3}$ = 0). We set the condition of $\alpha_{1}$ and $\alpha_{2}$ as 1.0. We use 6 sampled predictions to calculate the mean and standard deviation of metrics.
Training Metrics $\alpha_{3}=$ 0.0 $\alpha_{3}=$ 0.2 $\alpha_{3}=$ 0.4 $\alpha_{3}=$ 0.6 $\alpha_{3}=$ 0.8 $\alpha_{3}=$ 1.0
SFT
Words Composition
Semantic Similarity
Factuality
22.3 $\pm$ 6.5
110.4 $\pm$ 5.2
33.5 $\pm$ 66.1
23.2 $\pm$ 7.2
111.2 $\pm$ 4.8
36.8 $\pm$ 63.8
23.5 $\pm$ 7.5
$1$$\pm$ 5.1
35.8 $\pm$ 64.9
23.3 $\pm$ 7.7
110.9 $\pm$  5.9
37.2 $\pm$ 66.9
23.2 $\pm$ 8.1
110.8 $\pm$ 5.3
36.9 $\pm$ 65.3
24.2 $\pm$  6.8
$1$$\pm$ 5.1
38.2 $\pm$ 62.5
OLAPH (Step 1)
Words Composition
Semantic Similarity
Factuality
34.4 $\pm$ 7.1
112.1 $\pm$ 2.5
33.3 $\pm$ 67.2
35.3 $\pm$ 7.3
112.5 $\pm$ 2.1
51.2 $\pm$ 21.7
35.2 $\pm$ 8.2
112.6 $\pm$ 1.9
59.2 $\pm$ 18.8
36.1 $\pm$ 7.8
112.8 $\pm$ 2.0
66.3 $\pm$ 15.7
36.3 $\pm$ 7.5
113.1 $\pm$ 1.6
73.9 $\pm$ 13.8
36.9 $\pm$ 8.1
113.0 $\pm$ 1.4
81.2 $\pm$ 15.5
OLAPH (Step 2)
Words Composition
Semantic Similarity
Factuality
51.2 $\pm$ 4.9
112.0 $\pm$ 1.9
33.3 $\pm$ 68.1
52.3 $\pm$ 5.5
113.2 $\pm$ 1.8
54.1 $\pm$ 16.5
52.1 $\pm$ 5.6
112.9 $\pm$ 1.6
62.1 $\pm$ 13.2
52.3 $\pm$ 6.7
113.2 $\pm$ 1.5
68.8 $\pm$ 9.8
52.4 $\pm$ 6.7
112.9 $\pm$ 1.2
72.8 $\pm$ 8.5
55.7 $\pm$ 3.8
112.5 $\pm$ 0.9
86.5 $\pm$ 7.9
Table 6: Self-BioRAG 7B performance of sensitivity analysis ($\alpha_{3}$ = 0). We set the condition of $\alpha_{1}$ and $\alpha_{2}$ as 1.0. We use 6 sampled predictions to calculate the mean and standard deviation of metrics.
Training Metrics $\alpha_{3}=$ 0.0 $\alpha_{3}=$ 0.2 $\alpha_{3}=$ 0.4 $\alpha_{3}=$ 0.6 $\alpha_{3}=$ 0.8 $\alpha_{3}=$ 1.0
SFT
Words Composition
Semantic Similarity
Factuality
41.3 $\pm$ 12.2
121.1 $\pm$ 6.2
63.2 $\pm$ 28.9
40.5 $\pm$ 13.8
123.2 $\pm$ 5.8
64.5 $\pm$ 27.5
41.9 $\pm$ 11.9
123.5 $\pm$ 5.5
66.7 $\pm$ 25.5
42.5 $\pm$ 11.7
126.2 $\pm$ 7.2
69.9 $\pm$ 28.3
43.3 $\pm$ 9.9
125.9 $\pm$ 6.2
72.5 $\pm$ 24.9
43.2 $\pm$  10.1
125.0 $\pm$ 5.9
73.1 $\pm$ 25.8
OLAPH (Step 1)
Words Composition
Semantic Similarity
Factuality
53.8 $\pm$ 8.9
131.4 $\pm$ 4.9
61.1 $\pm$ 31.5
53.5 $\pm$ 9.0
132.2 $\pm$ 4.7
68.2 $\pm$ 18.9
53.3 $\pm$ 9.1
131.9 $\pm$ 3.5
73.2 $\pm$ 17.5
52.8 $\pm$ 8.0
131.3 $\pm$ 4.3
76.5 $\pm$ 13.3
52.6 $\pm$ 8.5
133.2 $\pm$ 3.9
87.3 $\pm$ 9.8
52.3 $\pm$ 7.5
135.7 $\pm$ 3.3
92.3 $\pm$ 11.2
OLAPH (Step 2)
Words Composition
Semantic Similarity
Factuality
54.3 $\pm$ 11.2
135.3 $\pm$ 3.8
62.3 $\pm$ 29.7
54.7 $\pm$ 7.9
135.2 $\pm$ 2.5
73.0 $\pm$ 15.9
53.2 $\pm$ 9.9
137.2 $\pm$ 2.3
77.2 $\pm$ 12.1
52.2 $\pm$ 9.1
136.9 $\pm$ 2.1
83.1 $\pm$ 9.9
54.1 $\pm$ 7.9
137.9 $\pm$ 1.8
91.2 $\pm$ 6.9
55.2 $\pm$ 8.3
138.2 $\pm$ 2.5
94.5 $\pm$ 8.9
Table 7: BioMistral 7B performance of using different thresholds to decide the preference and dispreference set. Threshold determines the quality and the size of the training dataset.
Training Metrics threshold = 0 threshold = 50 threshold = 100 threshold = 150 threshold = 200
SFT
Words Composition
Semantic Similarity
Factuality
Pairs of Dataset
21.8 $\pm$ 7.5
108.9 $\pm$ 4.7
33.8 $\pm$ 69.1
10,539
23.9 $\pm$ 6.9
110.1 $\pm$ 5.2
34.1 $\pm$ 63.5
7,532
23.2 $\pm$ 7.5
109.8 $\pm$ 3.5
33.3 $\pm$ 68.3
6,958
23.8 $\pm$ 7.1
109.2 $\pm$ 3.2
32.5 $\pm$ 71.3
6,472
24.2 $\pm$ 6.8
$1$$\pm$ 5.1
38.2 $\pm$ 62.5
5,173
OLAPH (Step 1)
Words Composition
Semantic Similarity
Factuality
Pairs of Dataset
33.9 $\pm$ 9.3
110.3 $\pm$ 4.5
66.2 $\pm$ 16.2
15,938
35.3 $\pm$ 9.9
112.1 $\pm$ 1.5
78.3 $\pm$ 12.3
13,521
35.5 $\pm$ 9.7
111.3 $\pm$ 2.2
75.8 $\pm$ 19.2
12,538
35.5 $\pm$ 9.7
110.5 $\pm$ 2.1
79.2 $\pm$ 17.3
10,529
36.9 $\pm$ 8.1
113.0 $\pm$ 1.4
81.2 $\pm$ 15.5
8,731
OLAPH (Step 2)
Words Composition
Semantic Similarity
Factuality
Pairs of Dataset
53.1 $\pm$ 4.2
111.2 $\pm$ 2.1
80.8 $\pm$ 9.1
20,331
54.2 $\pm$ 2.5
111.9 $\pm$ 0.8
85.7 $\pm$ 6.5
15,787
52.3 $\pm$ 4.3
113.9 $\pm$ 1.8
81.3 $\pm$ 7.2
3,029
53.2 $\pm$ 4.3
111.7 $\pm$ 1.2
84.5 $\pm$ 6.9
11,731
55.7 $\pm$ 3.8
112.5 $\pm$ 0.9
86.5 $\pm$ 7.9
10,832
A.3 Experimental Details